# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

### v1

@@ 5.3. The Synopsis Construction
In this subsection, we present a hierarchical tree synopsis, $Syn$, over auxiliary data for the dynamic graph $G_{t}$ (i.e., initial graph $G_{0}$ or its subsequent version at timestamp $t$) to accelerate online CD-SBN query processing.
@@
  • a keyword bit vector, $u.BV_{r}$, of user vertex $u$, which is a bit-OR of bit vectors $u_{i}.BV$ for all user vertices $u_{i}$ in $(2r)$-hop subgraph of $u$ (i.e., $u.BV_{r}=\bigvee_{\forall u_{i}\in hop(u,2r))}u_{i}.BV$);
  • a support upper bound, $u.ub\_sup_{r}$ of support upper bounds for edges in the $(2r)$-hop subgraph of $u$ (i.e.,$u.ub\_sup_{r}=max_{\forall e\in hop(u,2r)}{ub\_sup(e)}$), and;
@@
  • the maximum support upper bound $M_{i}.ub\_sup_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_sup_{r}}$, and;
  • the maximum score upper bound $M_{i}.ub\_score_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_score_{r}}$.
Bottom-up Construction of the Synopsis $\mathbf{Syn}$: We construct the hierarchical synopsis, $Syn$, as follows. For the initial bipartite graph $G_{0}$ (i.e., timestamp $t=0$), we start from each vertex $u_{i}\in U(G_{0})$ and perform a breadth-first search (BFS) to extract a $(2r)$-hop subgraph, $hop(u_{i},2r)$, centered at $u_{i}$ with radius $2r$ (for $r\in[1,r_{max}]$). Next, for each subgraph $hop(u_{i},2r)$, we compute its auxiliary data, $u_{i}.agg$. We sort the vertices in $G_{0}$ according to their summed score upper bounds for all possible radii (i.e., $\sum_{\forall r\in[1,r_{max}]}{u.ub\_score_{r}}$), and divide them into multiple partitions of the same size $\gamma$ (i.e., vertices with similar score upper bounds are grouped together), which are leaf nodes of the synopsis $Syn$ (as discussed in Section 5.3). Then, in a bottom-up manner, we recursively obtain non-leaf nodes $M$ of the synopsis $Syn$, by grouping leaf or non-leaf nodes $M_{i}$ (of the same size) on the lower level. The score upper bound associated with node $M$ is computed by aggregating the score upper bounds of its child nodes (i.e., $\sum_{\forall r\in[1,r_{max}]}$ $M_{i}.ub\_score_{r}$). The recursive synopsis construction terminates until one root node, $root(Syn)$, is obtained. After building the tree structure of the synopsis, we will recursively calculate aggregates, $M_{i}.agg$, of entries $M_{i}$ in leaf/non-leaf nodes of synopsis $Syn$. In addition, we keep an inverted list, $inv\_list$, for user vertices $u_{i}\in U(G_{t})$, where each user vertex $u_{i}$ is associated with an ordered list of node IDs from root to a leaf node in synopsis $Syn$ in which $u_{i}$ resides. This inverted list, $inv\_list$, can be used for accessing a particular user vertex for the synopsis maintenance (as discussed in Section 5.4).
@@ 5.4. Incremental Maintenance of the Synopsis
Specifically, if edge $e_{u_{i},v_{a}}$ is inserted or deleted, then we consider those users $u\in hop(u_{i},2r)$ within $(2r)$ hops away from $u_{i}$ as the affected vertices, which need to update keyword bit vectors $BV_{r}$. Moreover, since the support of each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$ may increase/decrease, user vertices, $u$, within $(2r)$-hop from $u_{i}$ may be affected (as $(2r)$-hop subgraph of $u$ may contain those edges $e$). Furthermore, the user relationship score between $u_{i}$ and each vertex $u_{j}\in N(v_{a})$ may change. Thus, we need to update the scores for those user vertices $u_{j}$ whose $(2r)$-hop subgraphs contain both $u_{i}$ and $u_{j}$.
Finally, in synopsis $Syn$, we identify those (non-)leaf nodes that contain the affected user vertices, whose aggregated data will be updated.
Maintenance of Leaf and Non-leaf Nodes: After identifying the affected vertices/nodes, we update their aggregates for all possible radii $r\in[1,r_{max}]$. Assume that an edge $e_{u_{i},v_{e}}$ is inserted or expired, and $u$ is an affected vertex. For the edge insertion, we merge $u_{i}.BV$ into $u.BV_{r}$, update $u.ub\_sup_{r}$ and $u.ub\_score_{r}$ for each affected vertex $u$, and update aggregates of the affected synopsis nodes $M$ containing $u$ (via $inv\_list$). For the edge expiration, we re-aggregate the data, $u.BV_{r}$, $u.ub\_sup_{r}$, and $u.ub\_score_{r}$, from $(2r)$-hop of each affected vertex $u$, and we re-aggregate the data in the affected nodes containing $u$ (if the re-aggregated data on the lower-level affected nodes have changed).
Maintenance of the Newly Added User Vertices: When an edge is inserted with a new user vertex $u_{new}$, we first compute the aggregates, $u_{new}.agg$ for this new user $u_{new}$ (within $(2r)$ hops from $u_{new}$) and add vertex $u_{new}$ to a leaf node $M_{i}$ in synopsis $Syn$ (containing a user vertex with the closest score upper bound). Next, we update the aggregate data of non-leaf nodes containing leaf node $M_{i}$.
[…]
@@ 6. CD-SBN query Processing
This section presents pruning strategies over synopsis $Syn$ to filter out unqualified synopsis nodes and provide algorithms for snapshot and continuous CD-SBN processing by accessing the synopsis.
@@ 6.1. Pruning via Synopsis
Keyword Pruning for Synopsis Entries: We first discuss how to use the aggregated keyword bit vector, $M_{i}.BV$, to rule out a synopsis entry $M_{i}$, if no query keyword in $Q$ exists in the vertices under $M_{i}$.
@@
(Synopsis Keyword Pruning) Given a synopsis entry $M_{i}$ and a set, $Q$, of query keywords, entry $M_{i}$ can be safely pruned, if it holds $M_{i}.BV\bigwedge Q.BV=\boldsymbol{0}$, where $Q.BV$ is a bit vector hashed from query keywords in $Q$.
@@
Since $M_{i}.BV$ is an aggregated keyword bit vector computed by $\bigvee_{\forall u\in M_{i}}u.BV$, the $f(w)$-th bit position of $M_{i}.BV$ with the value 00 means that the keyword $w$ is not contained in the keyword set, $v.K$, of any item vertex $v$, where $v\in N(u)$ and the $f(\cdot)$ is a hash function. Therefore, if $M_{i}.BV\wedge Q.BV=\boldsymbol{0}$ holds, no query keywords in $Q$ belong to $M_{i}$, which violates the keyword constraint in Definition 2.8 (i.e., $v.K\cap Q\neq\emptyset$). In this case, the synopsis entry $M_{i}$ cannot contain any vertices in the CD-SBN query answer and thus can be safely pruned, which completes the proof. ∎
@@ 6.2. Snapshot CD-SBN Processing Algorithm
Input: i) a streaming bipartite network $G_{t}$; ii) a tree synopsis $Syn$ over $G_{t}$, and; iii) a snapshot CD-SBN query with a set, $Q$, of query keywords, a query support threshold $k$, a query maximum radius $r$, and a threshold, $\sigma$, of user relationship score
@@ 10       $(M,key)$ = $(\mathcal{H})$;
Initialization: Given a query keyword set $Q$, the algorithm first hashes all query keywords in $Q$ into a query keyword bit vector $Q.BV$ (line 1). Then, we initialize a maximum heap $\mathcal{H}$ (for synopsis traversal), accepting heap entries in the form of $(M,key)$, where $M$ is a synopsis entry and $key$ is the maximum score upper bound, $M.ub\_score_{r}$, of all user vertices in $M$ (as mentioned in Section 5.3) (line 2). Intuitively, a heap entry with a large key (i.e., large maximum score upper bound) is more likely to contain communities with high user relationship scores. Next, we insert the root of the tree synopsis $Syn$, in the form $(root(Syn),+\infty)$, into the heap $\mathcal{H}$ (line 3). Moreover, we use an initial empty set $P$ to keep the candidate subgraphs and an empty set $R$ to store snapshot CD-SBN query answers (line 4).
Synopsis Traversal: Next, we utilize the maximum heap $\mathcal{H}$ to traverse the synopsis (lines 5-18). Each time we pop out a heap entry $(M,key)$ from heap $\mathcal{H}$ with the highest key, $key$ (i.e., the maximum score upper bound) (lines 5-6). If $key$ is less than the score threshold, $\sigma$, it indicates that all the remaining heap entries in $\mathcal{H}$ have a score less than $\sigma$, and they cannot be contained in our CD-SBN query results. Thus, we can safely terminate the synopsis traversal early (lines 7-8).
When $M$ is a leaf node, we consider each user vertex $u_{i}\in M$ and check whether or not we can find any candidate community centered on $u_{i}$ (lines 9-14). For each user $u_{i}\in M$, if the $2r$-hop subgraph $hop(u_{i},2r)$ cannot be pruned by our pruning strategies (i.e., keyword, support, and score upper bound pruning), then we will obtain a maximal $(k,r,\sigma)$-bitruss $g$ (satisfying the keyword constraint) within $hop(u_{i},2r)$ (lines 10-12). If such a subgraph $g$ exists, we will add $g$ to the candidate community set $P$ (lines 13-14).
On the other hand, when $M$ is a non-leaf node, we will check each entry $M_{i}$ of node $M$ (lines 15-16). If synopsis entries $M_{i}$ cannot be pruned using synopsis-level pruning strategies, then we will insert $(M_{i},M_{i}.ub\_score_{r})$ into the heap $\mathcal{H}$ for further checking (lines 17-18).
Refinement: Finally, we remove duplicate subgraphs from $P$ (due to overlap of $2r$-hop subgraphs), refine candidate communities in $P$ to obtain actual CD-SBN community answers in $R$, and return $R$ (lines 19-20).
@@
6.3. Continuous CD-SBN Processing Algorithm
Algorithm 3 details our continuous CD-SBN query processing approach to monitoring answers. First, we initialize the CD-SBN result set $R_{t}$ at timestamp $t$ with that $R_{t-1}$ at timestamp $(t-1)$ (line 1). Next, we check the constraints of the existing communities to refine the previous result set (lines 2-6). Then, we obtain the potential communities from the $2r$-hop of user vertices near the insertion edge and remove the duplication (lines 7-10) Finally, we merge the candidate communities into the result set and return the result set at timestamp $t$ (lines 11-12).
Input: i) a streaming bipartite network $G_{t-1}$ with an update stream $S$; ii) a set, $R_{t-1}$, of community answers at timestamp $(t-1)$, and; iii) a continuous CD-SBN query with a set, $Q$, of query keywords, a support threshold $k$, a maximum radius $r$, and a threshold, $\sigma$, of user relationship score
@@ 6.3. Continuous CD-SBN Processing Algorithm
// Processing the Expired Item $p_{t-s}=(e_{u_{i},v_{a}},t-s)\in S$
4
5for each community answer $g\in R_{t}$ do

### v2

@@ 5.3. The Synopsis Construction
We next present a hierarchical tree synopsis, $Syn$, on auxiliary data for dynamic graph $G_{t}$ (i.e., initial graph $G_{0}$ or its subsequent version at timestamp $t$) to accelerate online CD-SBN query processing.
@@
  • a keyword bit vector, $u.BV_{r}$, of user vertex $u$, which is a bit-OR of bit vectors $v_{i}.BV$ for all item vertices $v_{i}$ in $(2r)$-hop subgraph of $u$ (i.e., $u.BV_{r}=\bigvee_{\forall v_{i}\in hop(u,2r))}v_{i}.BV$);
[…]
@@
  • the maximum support upper bound $M_{i}.ub\_sup_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_sup_{r}}$;
  • the maximum score upper bound $M_{i}.ub\_score_{r}=\\ max_{\forall u\in M_{i}}{u.ub\_score_{r}}$, and;
  • a pointer, $M_{i}.ptr$, pointing to a child node.
Bottom-up Construction of the Synopsis $\mathbf{Syn}$: We construct the hierarchical synopsis, $Syn$, as follows. For the initial bipartite graph $G_{0}$ (i.e., timestamp $t=0$), we start from each vertex $u_{i}\in U(G_{0})$ and perform a breadth-first search (BFS) to extract a $(2r)$-hop subgraph, $hop(u_{i},2r)$, centered at $u_{i}$ with radius $2r$ (for $r\in[1,r_{max}]$). Next, for each subgraph $hop(u_{i},2r)$, we compute its auxiliary data, $u_{i}.agg$. We sort the vertices in $G_{0}$ according to their summed score upper bounds for all possible radii (i.e., $\sum_{\forall r\in[1,r_{max}]}{u.ub\_score_{r}}$), and divide them into multiple partitions of the same size $\gamma$ (i.e., vertices with similar score upper bounds are grouped together), which are leaf nodes of the synopsis $Syn$ (as discussed in Section 5.3). Then, in a bottom-up manner, we recursively obtain non-leaf nodes $M$ of the synopsis $Syn$, by grouping leaf or non-leaf nodes $M_{i}$ (of the same size) on the lower level. The score upper bound associated with node $M$ is computed by aggregating the score upper bounds of its child nodes (i.e., $\sum_{\forall r\in[1,r_{max}]}$ $M_{i}.ub\_score_{r}$). The recursive synopsis construction terminates until one root node, $root(Syn)$, is obtained. After building the tree structure of the synopsis, we will recursively calculate aggregates, $M_{i}.agg$, of entries $M_{i}$ in leaf/non-leaf nodes of synopsis $Syn$. In addition, we keep an inverted list, inv_listinv\_list, for user vertices $u_{i}\in U(G_{t})$, where each user vertex $u_{i}$ is associated with an ordered list of node IDs from root to a leaf node in synopsis $Syn$ in which $u_{i}$ resides. This inverted list, inv_listinv\_list, can be used for accessing a particular user vertex for the synopsis maintenance (as discussed in Section 5.4).
Complexity Analysis: For a tree index $\mathcal{I}$, we denote the fanout of each non-leaf node $N$ as $\gamma$ and the average number of vertex degrees as avg_degavg\_deg. To compute the auxiliary data that we mentioned above, it takes $O(2r_{max})$ to aggregate the item keyword bit vector ($v_{i}.BV_{r}$), support upper bound ($ub\_sup_{r}$), and user relationship score upper bound $ub\_score_{r}$ from child nodes. Since the number of leaf nodes is equal to the number of user vertices $|U(G)|$ divided by $\gamma$, the depth of the tree index $\mathcal{I}$ is $\lceil\log_{\gamma}{|U(G)|}\rceil$. The time complexity of constructing a $\gamma$-ary tree with a depth of depdep is $O(\gamma^{dep}-1)/gamma-1$. Therefore, the total time complexity of our tree index construction is given by $O(2r_{max}\cdot(\gamma^{\lceil\log_{\gamma}{|U(G)|}\rceil}-1)/(\gamma-1))$.
For the space complexity, each leaf node contains a keyword bit vector of size $B$, a support upper bound, and a user relationship score upper bound. Therefore, the space complexity of all leaf nodes is $O(2r_{max}\cdot B\cdot|V(G)|/\gamma)$. On the other hand, for each non-leaf node, it needs a bit vector sized of $B$, a support upper bound, a user score upper bound, and a list of pointers with size $\gamma$ to its child nodes. Since the depth of the tree index is $\lceil\log_{\gamma}{|V(G)|}\rceil$, the space complexity of all non-leaf nodes is $O(2r_{max}\cdot(B+\gamma)\cdot(\gamma^{\lceil\log_{\gamma}{|V(G)|}\rceil-1}-1)/(\gamma-1))$
Overall, the space cost of our tree index is $O(2r_{max}\cdot((B+\gamma)\cdot(\gamma^{\lceil\log_{\gamma}{|V(G)|}\rceil-1}-1/(\gamma-1))+B\cdot|V(G)|/\gamma)))$.
@@ 5.4. Incremental Maintenance of the Synopsis
Specifically, if the weight of edge $e_{u_{i},v_{a}}$ is updated, we will consider those users $u\in hop(u_{i},2r)$ within $(2r)$ hops away from $u_{i}$ as the affected vertices, which need to update keyword bit vectors $BV_{r}$. Moreover, since the support of each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$ may increase/decrease, user vertices, $u$, within $(2r)$-hop from $u_{i}$ may be affected (as $(2r)$-hop subgraph of $u$ may contain those edges $e$). Moreover, the user relationship score between $u_{i}$ and each vertex $u_{j}\in N(v_{a})$ may change. Thus, we need to update the scores of user vertices $u_{j}$ whose $(2r)$-hop subgraphs contain both $u_{i}$ and $u_{j}$.
Finally, in synopsis $Syn$, we identify those (non-)leaf nodes that contain the affected user vertices and update their aggregated data.
Maintenance of Leaf and Non-leaf Nodes: After identifying the affected vertices/nodes, we update their aggregates for all possible radii $r\in[1,r_{max}]$. Assume that an edge $e_{u_{i},v_{a}}$ is inserted/expired, and $u$ is an affected vertex. For the edge insertion, we merge $v_{a}.BV$ into $u.BV_{r}$, update $u.ub\_sup_{r}$ and $u.ub\_score_{r}$ for each affected vertex $u$, and update aggregates of the affected synopsis nodes $M$ containing $u$ (via inv_listinv\_list). For the edge expiration, we re-aggregate the data, $BV_{r}, u.u$$ub\_sup_{r}$, and $u.ub\_score_{r}$, from $(2r)$-hop of each affected vertex $u$, and we re-aggregate the data in the affected nodes containing $u$ (if the re-aggregated data on the lower-level affected nodes have changed).
Maintenance of the Newly Added User Vertices: When an edge is inserted with a new user vertex $u_{new}$, we first compute the aggregates, $u_{new}.agg$ for this new user $u_{new}$ (within $(2r)$ hops from $u_{new}$) and add $u_{new}$ to a leaf node $M_{i}$ in synopsis $Syn$ (containing a user vertex with the closest score upper bound). Next, we update the aggregate data of non-leaf nodes containing leaf node $M_{i}$.
Time Complexity Analysis: Assume that an edge $e_{u_{i},v_{a}}$ is inserted/expired. To identify affected vertices, it takes $O(|U(hop(u_{i},2r)|)$ to obtain vertices that may need to update $BV_{r}$, $u.ub\_sup_{r}$, and $u.ub\_score_{r}$. After that, we check inv_listinv\_list to retrieve the affected nodes containing the affected vertices with the time complexity of $O(1)$. When edge $e_{u_{i},v_{a}}$ is inserted, it takes $O(1)$ to merge $v_{a}.BV$ into $u.BV_{r}$, maintain support upper bound $u.ub\_sup_{r}$ and score upper bound $u.ub\_score_{r}$ for each affected vertex $u$. Then, we update the aggregates in leaf and non-leaf nodes containing $u$ with a cost of $O(1)$. For the edge expiration, it takes $O(|U(hop(u,$ $2r))|)$ time complexity to aggregate the bit vector of user keywords, calculate support upper bound $u.ub\_sup_{r}$, and update score upper bound $u.ub\_score_{r}$ according to the $(2r)$-hop subgraph of each affected vertex $u$. To maintain non-leaf nodes, we recompute the aggregates for each non-leaf node, which takes $O(\gamma)$, where $\gamma$ is the average size of synopsis nodes. In the worst case, we can find $|U(hop(u_{i},2r))|$ affected vertices and $h$ synopsis nodes for each vertex, where $h$ is the height of the tree synopsis $Syn$. In summary, it takes $O(|U(hop(u_{i},2r))|+h)$ for each edge insertion and $O(|U(hop(u_{i},2r))|\cdot|U(hop(u,2r))|+h\cdot\gamma)$ for each edge expiration.
@@ 6.1. Pruning via Synopsis
Keyword Pruning for Synopsis Entries: We first discuss how to use the aggregated keyword bit vector, $M_{i}.BV_{r}$, to rule out a synopsis entry $M_{i}$, if no query keyword in $Q$ exists in the vertices under $M_{i}$.
@@
(Synopsis Keyword Pruning) Given a synopsis entry $M_{i}$ and a set, $Q$, of query keywords, entry $M_{i}$ can be safely pruned, if it holds $M_{i}.BV_{r}\bigwedge Q.BV=\boldsymbol{0}$, where $Q.BV$ is a bit vector hashed from query keywords in $Q$.
@@
Since $M_{i}.BV$ is an aggregated keyword bit vector computed by $\bigvee_{\forall v\in M_{i}}v.BV$, the $f(w)$-th bit position of $M_{i}.BV_{r}$ with the value 0 means that the keyword $w$ is not contained in the keyword set, $v.K$, of any item vertex $v$, where $v\in N(u)$ and the $f(\cdot)$ is a hash function. Therefore, if $M_{i}.BV_{r}\wedge Q.BV=\boldsymbol{0}$ holds, no query keywords in $Q$ belong to $M_{i}$, which violates the keyword constraint in Definition 2.8 (i.e., $v.K\cap Q\neq\emptyset$). In this case, the synopsis entry $M_{i}$ cannot contain any vertices in the CD-SBN query answer, and thus it can be safely pruned, completing the proof. ∎
@@ 6.2. Snapshot CD-SBN Processing Algorithm
Input: i) a streaming bipartite network $G_{t}$; ii) a tree synopsis SynSyn over $G_{t}$, and; iii) a snapshot CD-SBN query with a set, $Q$, of query keywords, a query support threshold $k$, a query maximum radius $r$, and a user relationship score threshold $\sigma$
@@ 10  $(M,key)$ = de-heap()($\mathcal{H}$);
Initialization: Given a query keyword set $Q$, the algorithm first hashes all query keywords in $Q$ into a query keyword bit vector $Q.BV$ (line 1). Then, we initialize a maximum heap $\mathcal{H}$ (for synopsis traversal), accepting heap entries in the form of $(M,key)$, where $M$ is a synopsis entry and keykey is the maximum score upper bound, $M.ub\_score_{r}$, of all user vertices in $M$ (as mentioned in Section 5.3) (line 2). Intuitively, a heap entry with a large key (i.e., a large maximum score upper bound) is more likely to contain communities with high user relationship scores. Next, we insert the root of the tree synopsis $Syn$, in the form $(root(Syn),+\infty)$, into the heap $\mathcal{H}$ (line 3). Moreover, we use an initial empty set $P$ to keep the candidate subgraphs and an empty set $R$ to store snapshot CD-SBN query answers (line 4).
Synopsis Traversal: Next, we utilize the maximum heap $\mathcal{H}$ to traverse the synopsis (lines 5-18). Each time we pop out a heap entry $(M,key)$ from heap $\mathcal{H}$ with the highest key, keykey (i.e., the maximum score upper bound) (lines 5-6). If keykey is less than the score threshold, $\sigma$, it indicates that all the remaining heap entries in $\mathcal{H}$ have a score less than $\sigma$, and they cannot be contained in our CD-SBN query results. Thus, we can safely terminate the synopsis traversal early (lines 7-8).
When $M$ is a leaf node, we consider each user vertex $u_{i}\in M$ and check whether a candidate community centered on $u_{i}$ exists (lines 9-14). For each user $u_{i}\in M$, if the $2r$-hop subgraph $hop(u_{i},2r)$ cannot be pruned by our pruning strategies (i.e., keyword, support, and score upper bound pruning), then we will obtain a maximal $(k,r,\sigma)$-bitruss $g$ (satisfying the keyword constraint) within $hop(u_{i},2r)$ (lines 10-12). If such a subgraph $g$ exists, we will add $g$ to the candidate community set $P$ (lines 13-14).
On the other hand, when $M$ is a non-leaf node, we will check each entry $M_{i}$ of node $M$ (lines 15-16). If synopsis entries $M_{i}$ cannot be pruned by synopsis-level pruning strategies, then we will insert $(M_{i},M_{i}.ub\_score_{r})$ into the heap $\mathcal{H}$ for next checking (lines 17-18).
Refinement: Finally, we remove duplicate subgraphs from $P$ (due to overlap of $2r$-hop subgraphs), refine candidate communities in $P$ to obtain CD-SBN query answers in $R$, and return $R$ (lines 19-20).
@@
Input: i) a streaming bipartite network $G_{t-1}$ with an update stream $S$; ii) a set, $R_{t-1}$, of community answers at timestamp (t1)(t-1), and; iii) a continuous CD-SBN query with a set, $Q$, of query keywords, a support threshold $k$, a maximum radius $r$, and a threshold, $\sigma$, of user relationship score
@@ 1 $R_{t}=R_{t-1}$;
4for each community answer $g\in R_{t}$ do